OpenStudy (anonymous):

5x + 2y = 8 x + y = 4 If you want to solve the system of equations by addition, which of the following could you do? Multiply the second equation by 5 and add. Multiply the second equation by 2 and add. Multiply the second equation by -2 and add.

OpenStudy (trojanpoem):

third choice.

OpenStudy (trojanpoem):

5 x + 2y = 8 -2x - 2y = -8 x = 0 y = 4

OpenStudy (anonymous):

ok thanks:)